# S2Edit: Text-Guided Image Editing with Precise Semantic and Spatial Control

### v1

@@ 4.4 Ablation Analysis
Effect of Each Component. In Fig. 8, we analyze the effect of all three components within S2Edit: Identity Fine-tuning (IFT), Semantic Control (SeC), and Spatial Control (SpC). Without them, our method falls back to the Null-text Inversion (Mokady et al., 2023) baseline, where the person’s identity information is completely lost in the edited image. Although this issue can be resolved by naively applying Identity Fine-tuning to learn an identity token [I], which fundamentally aligns with DreamBooth (Ruiz et al., 2023) that preserves the person’s appearance, the target attribute included in the prompt (“bangs”) is missing in the edited image. The attention map of [I] reveals that this is caused by excessive information being encoded in [I], which limits the editability of the person’s hair. Adding Semantic Control helps alleviate this problem by disentangling [I] from information contained in the original prompt. However, the lack of spatial constraints on the attention map causes [I]’s attention to drift towards irrelevant regions such as the background, while focusing less on the person’s face. The generated image under Identity Fine-tuning and Semantic Control aligns with the target prompt well, yet it distorts subtle details like the person’s skin tone. Finally, by adding Spatial Control that exclusively directs [I]’s attention to the region of interest, the edited image from our full method not only accurately follows the target prompt to add bangs, but also keeps the person’s identity intact.
@@ 5 Conclusion
Limitations. A text prompt describing the source image is required for editing, which is a common constraint in methods building on pre-trained text-to-image diffusion models (Hertz et al., 2022; Zhang et al., 2023). A future direction is thus to enable editing without source prompts, which may be achieved by inverting prompts from images (Dong et al., 2023) or captioning models (Li et al., 2022).

### v2

@@ 4.4 Ablation Analysis
Effect of Each Component. In Fig. 8, we analyze the effect of all three components within S2Edit: Identity Fine-tuning (IFT), Semantic Control (SeC), and Spatial Control (SpC). Without them, our method falls back to the Null-text Inversion (mokady2023null) baseline, where the person’s identity information is completely lost in the edited image. Although this issue can be resolved by naively applying Identity Fine-tuning to learn an identity token [I], which fundamentally aligns with DreamBooth (ruiz2023dreambooth) that preserves the person’s appearance, the target attribute included in the prompt (“bangs”) is missing in the edited image. The attention map of [I] reveals that this is caused by excessive information being encoded in [I], which limits the editability of the person’s hair. Adding Semantic Control helps alleviate this problem by disentangling [I] from information contained in the original prompt. However, the lack of spatial constraints on the attention map causes [I]’s attention to drift towards irrelevant regions such as the background, while focusing less on the person’s face. The generated image under Identity Fine-tuning and Semantic Control aligns with the target prompt well, yet it distorts subtle details like the person’s skin tone. Finally, by adding Spatial Control that exclusively directs [I]’s attention to the region of interest, the edited image from our full method not only accurately follows the target prompt to add bangs, but also keeps the person’s identity intact.
@@ 5 Conclusion
Limitations. A text prompt describing the source image is required for editing, which is a common constraint in methods building on pre-trained text-to-image diffusion models (hertz2022prompt; SINE). A future direction is thus to enable editing without source prompts, which may be achieved by inverting prompts from images (PromptTuningInversion) or captioning models (li2022blip).